OpenStudy (anonymous):

Solving by substitution! Help, will medal. I have 2 equations that need to be solved by substitution and I'm having trouble seeing what I did wrong. -4.5x-2y=-12.5 3.25x-y=-.75 I got (0,4) and my teacher said it's supposed to be (1,4). can you show work so i know what i did wrong?

OpenStudy (tkhunny):

Well, you should show your work. Let's see where youwandered off.

OpenStudy (anonymous):

I picked the first equation, -4.5x-2y=-12.5 I added 4.5x to both sides. -2y=4.5x-12.5 /-2 /-2 /-2 y=-2.25x+6.25 I replaced it back into the original equation. -4.5x-2(-2.25x+6.25)=-12.5 -4.5x+4.5x-12.5=-12.5 the two variables cancel each other out and so does -12.5 when you transfer it to the other side of the "=".

OpenStudy (tkhunny):

So close!! Right idea. Don't put it back into the equation where it came from. That will always result in something unhelpful. Put it in the OTHER equation! 3.25x-(-2.25x+6.25)=-.75 Now, give it a go.